OpenStudy (anonymous):

How do you solve x-2y=1 y=-x+4 by using substitution?

OpenStudy (anonymous):

Well, you know what y is:, so put y in the second equation: x - 2(-x+4) = 1 x + 2x - 8 = 1 3x -8 = 1 3x = 9 x = 3 Then, plug for x in the original equation to get: y = -(3) + 4 = 1 Therefore, x = 3 and y = 1.

OpenStudy (anonymous):

Thank you sir!